# Using machine learning to measure evidence of students’ sensemaking in physics courses

### v2

@@ II.2 Post-Hoc Measurement as Multi-Label Binary Annotation
Step 1. Select problems from a particular subject area (e.g. classical mechanics), where there exists a physical phenomena to be sensemade in each problem as a part of the strategy to solution.
@@
Step 2. For each problem and for each domain , determine the set of yes/no criteria questions that are pertinent to the specific problem. For example, in Problem 4 (II.1), there is only a single object and thus with the criteria question Does the student identify the object?. However, there are two positioning criteria questions that are relevant to sensemaking an object’s positioning under a net external influence: Does the student identify the object positioning before the net applied force? Does the student identify the object positioning after the net applied force? The annotator should end up with the total number of criteria questions per problem as .
@@
Step 3. Given a student’s text response to problem , the annotator will use the yes/no criteria questions to obtain a vector of binary digits representing their answers. The resulting criteria vector for a student response is denoted as . To further compute a sensemaking score one can sum the number of and divide by the length of the vector . This score is to be interpreted as the amount of evidence of physical sensemaking. An annotator can also decide to weight the domains (or specific criteria) differently. In this work, we consider all domains and criteria to equally contribute to the sensemaking score. To finish off our example, the criteria vector for the student explanation in Fig. 1 is with the corresponding sensemaking score . The student achieved such a high score by imagining a scenario where multiple interactions are influencing the boxes’ dynamics and communicating the resulting mechanistic relationship between a net external force and an the boxes’ change in motion. The student received a yes for all criteria except for the one in the property domain. The full list of criteria for Problem 4 is displayed in the Appendix A).
II.3 Theoretical Benefits and Limitations
Utility. This annotation scheme is intentionally designed to measure evidence of physical sensemaking given student explanations. The measurements allow for evidence of sensemaking to be quantitatively compared to other numerical values of interest, and distinguished from correctness. For example, evidence of sensemaking is different from measuring sensemaking correctness (i.e. the student is engaging in the sensemaking process with information that is determined by the authority figure to be canonically correct). The current scheme could be easily modified for this other purpose; the annotator would simply need to choose to annotate for both sensemaking and correctness according to the problem (e.g. the student identifies the correct positioning). However in this work, we stick to sensemaking scores that measure evidence of sensemaking, and leave correctness for future studies.
Moreover, we emphasize that low sensemaking scores measured do not indicate that students cannot sensemake. It simply measures that in the current snapshot, evidence of sensemaking is low. There could be several reasons as to why the evidence is low (i.e. the problem itself didn’t encourage sensemaking, the student applied a memorized fact or procedure, etc.). Put concretely, our annotation scheme does not allow one to make any claims about causality.
Flexibility. Individuals can design or select their own problem sets, and then annotate the student explanations according to the same 8 domains. In proposing such an adaptive scheme, we attempt to balance imposing a systematic procedure with allowing the annotator to have enough freedom to make decisions.
Reproducibility. We provide a step-by-step outline of the annotation procedure, concise definitions of the 8 domains, and annotated examples of 385 student explanations across 4 physics problems of different criteria vector lengths. Our hope is that this content leaves little ambiguity for a new annotator looking to reproduce our results or implement the scheme in their own work. However, we acknowledge that some variability among annotators could exist and be quantified in the future with inter-rater reliability studies [39].
Reducibility. Each criteria vector length is the amount of binary information (bits) required to represent the evidence of a student’s sensemaking with respect to a particular problem. It is indeed possible that the 8 domains and criteria could be reduced such that a fewer number of binary variables are needed. While a reduced model form is not currently known; we speculate that such a form could exist.
Scalability. We acknowledge that manual annotation of a large number of student responses would require many hours of labor, often infeasible in practical settings. For example, the student explanation in Fig 1 took minutes to annotate. For a class of one-hundred students, this extrapolates to annotation hours with respect to a single problem. To resolve this scalability challenge, we show how the procedure can be automated by training an ML classification model on student explanations.
II.4 Sensemaking Related Works
Odden and Russ explicitly define sensemaking as “the dynamic process of building an explanation to resolve a perceived gap or conflict in knowledge.” The authors provide examples of individuals engaging in this process, and discuss how sensemaking can be conceptualized in three ways: a stance or framing (i.e. an individuals’ expectation of themselves in their current situation), a cognitive process built on the theory of knowledge integration, and a discourse practice related to argumentation. Each conception addresses what sensemaking is with respect to a different system level of study. Cognitive theories are based on the mental components and interactions happening during sensemaking, whereas framing can be considered one hierarchal level up as a latent description of a student’s current expectations in their environment when sensemaking is taking place. Lastly, discourse analysis is based on studying the content within and form of conversational dialogue between students.
Qualitative analysis has shaped our understanding of how sensemaking begins, functions, and ends for students. Odden and Russ [15] demonstrate how vexing questions stimulate the sensemaking process and can sustain it for longer durations of time than typically observed. Benedict-Chambers et al. [40] provides four specific types of questions focused on guiding students towards sensemaking including explication, explanation, science concept, and science practice questions. To characterize the explicit features of in-moment sensemaking, Odden and Russ [23] propose that the process be modeled as an epistemic game. This model introduces sensemaking as a three step procedure: 1. noticing a gap or inconsistency, 2. generating a mechanistic explanation, and 3. resolving the gap or inconsistency. The authors discuss a precursor to sensemaking, which includes assembling a knowledge framework by initially discussing ideas socially. This brainstorming step is necessary for one to eventually recognize a gap in their own knowledge system. During the game, the students can make moves (e.g. refining a question, appealing to reality, building an analogy, etc.) and if certain constraints are violated, students exit the game play. Three insightful constraints are discussed: 1. sensemaking is a discussion between individuals rather than an imbalanced monologue, 2. sensemaking occurs when there is a phenomena to sensemake that has not already been sensemade, and 3. sensemaking contains self-generated definitions of ideas alongside formal definitions. Kapon [20] puts forth a metric for how explanations generated during the sensemaking process are implicitly evaluated by individuals - based on intuitive knowledge, mechanism, and framing. More specifically, students evaluate explanations as “good enough” based on whether it aligns with their phenomenological experiences, is mechanistic rather than teleological, and satisfies their immediate expectations of their own understanding. Odden [19] explores how conceptual blends support sensemaking through a qualitative case study of students using concepts from kinematics to understand the mechanisms that underlie the phenomena in electric circuits. The author proposes the approach of generalizing conceptual ideas from one sub-area of physics to another as a mechanism that allows students to make connections in the sensemaking process. Furthermore, Sirnookar et al. [18] explores the role of modeling in sensemaking and analyzes the interactions of these two processes as two students are interviewed while solving a physics problem. The authors find that barriers in modeling can inhibit students’ sustained sensemaking.
In pursuit of the longer-term impacts of sensemaking, Radoff et al. [14] provides a qualitative account of a student engaging in sensemaking throughout an academic semester. Interviews and coursework demonstrated that experimentation with sensemaking caused more positive associations with learning, taking on challenges, and constructive confusion. In Kuo et al. [31], the authors show through a controlled classroom experiment that students who are instructed with mathematical sensemaking stimulation are prone to obtaining solutions to physics problems that are more accurate and insightful than those who did not receive this type of explicit instruction.
The sensemaking literature is fragmented across several scientific disciplines and is also typically broken into categories of scientific and mathematical sensemaking, where more recent work has attempted to define a blended model of the two [16, 26, 27]. Kaldaras and Wieman [27] build upon prior works to put forth a cognitive model of blended math-science sensemaking that introduces three levels of sophistication: 1. developing qualitative relationships among relevant variables in mathematical equations that describe a particular phenomenon, 2. developing mathematical relationships among these variables, and 3. explaining how the mathematical operations used in the equation connect back to the physical phenomena. The authors attempt to validate their model by analyzing student responses from interviews, specifically probing the various levels of sophistication in the framework. While the authors found convincing evidence that the levels of sophistication were present across all student responses, it remained inconclusive as to whether there exists an actual sophistication ranking in the levels. As a result, it is still uncertain whether there exists a hierarchy among cognitive connections in math-science sensemaking. This work is directly linked with research that asks: what does it mean to understand a physics equation? [34] and seeks out structure in symbolic expressions that aid in student understanding.
The works above collectively agree on a key characteristic of the sensemaking process: the dynamic generation of mechanistic explanations for a given phenomena. When students have resolved their inconsistency, they have settled on a mechanistic explanation. The emphasis on mechanism in sensemaking is historically tied with research that characterizes student’s mechanistic reasoning in the classroom [17, 25, 41]. Russ et al. is arguably the the most influential contribution made thus far in defining mechanism in the context of STEM Ed., providing a framework for discourse analysis developed from ideas in philosophy of science [25]. In the framework, mechanistic reasoning is characterized as a process that uses experience-based evidence to construct causal relationships among objects describing “how underlying structure and activities can account for observable changes in a system”. The work puts forth a coding scheme of nine categories: 1. describing the target phenomenon they are trying to explain, 2. identifying setup conditions of the spatial and temporal organization of entities that construct the phenomena, 3. identifying the entities as components that construct the phenomena, 4. identifying activities (actions and interactions) of the entities, 5. identifying properties of entities, 6. identifying the spatial organization of entities relative to each other, 7. chaining reasoning strategy for backward or forward inference, 8. analogies, and 9. mental model animation. This framework has been widely adopted by the STEM Ed. community for qualitatively analyzing student reasoning, and we highlight that it overlaps significantly with our 8 domains presented to measure the evidence of sensemaking. There are a few key differences. The first is that we separate entities into objects and influences, where influences interact with objects in such a way that only the influence object effect is relevant to the particular phenomena. This is especially useful when dealing with fields in physics, which are impacted from object interactions, but the effects are oftentimes ignored in classical settings or introductory courses. Our two domains positioning and movement track both the set-up and subsequent conditions of the dynamics of each object. The positioning domain also includes how the objects relate to one another structurally (i.e. object A is on top of object B) when this is relevant for the particular phenomena. As such, we reduce three of the codes (2,6,9) in Russ et al. to a single domain. As we aim to measure post-hoc explanations rather than in-moment discourse, we can code for the chaining and analogy creation process (which can be viewed as sub-sensemaking processes) proposed in Russ et al. by annotating whether the student identifies the relevant mechanistic relationships. In addition, we add a domain for descriptive relationships, which are the relations between the properties of objects - more akin to correlation than causal mechanism. Otherwise, with respect to Russ et al., our properties domain directly maps to code 5 and our interactions domain directly maps to code 4. As previously stated, our work builds off of the numerous qualitative studies discussed above that have identified the features of sensemaking and mechanistic reasoning through analyzing student discourse. We view our main contribution as transforming these features into 8 domains that contain sub-questions with binary outcomes such that we can obtain a post-hoc sensemaking measurement of a student’s explanation with respect to any particular physics problem. Introducing such a binary encoding scheme allows one to quantify evidence of student sensemaking and naturally lends itself to automating the measurement scheme with ML.
III How to Automate Measurements
III.1 Multi-Label Binary Classification Model
ML models have been shown to be capable of replacing labor-intensive encoding of qualitative data with automated procedures [41, 42, 43, 44, 45]. In this work, we only consider models which are open-access, which allows us to train classifiers for our specific task and does not require us to share student data with a 3rd party. Furthermore, we limit our consideration to models which are capable of running on consumer hardware. In our scenario, the task is the same as the human annotator’s: to map student text explanations across independent problems to sensemaking scores .
@@ III.1 Multi-Label Binary Classification Model
From text to embeddings. Given a student explanation , we can select an encoder to map all text in the response to a numerical representation . A simple encoder design is counting; we determine a vocabulary as the set of words across all student explanations in the training dataset, with an additional vocabulary entry for unknown words, and record the number of times each one appears in the example. This numerical representation, or embedding , typically called BagofWords, reduces the richness of the text explanation to the pattern of word occurrence [46, 47]. From here on, we will refer to this encoder as BoW.
@@
To capture more complex patterns (e.g relative co-occurrence between all words in the explanation, in-context word occurrence, and word positional associations with grammar rules), one can use a large language model (LLM) encoder [48]. For example, Google’s BERT (Bi-directional Encoder Representations from Transformers) model combines tokenization, self-attention, and a multi-layer perceptron neural network - all pre-trained on English Wikipedia and BookCorpus to generate token embeddings [35]. A token is a set of characters containing a unique ID in BERT’s vocabulary, such that an entire text can be converted to tokens with individual embeddings , which carry information about all other tokens in the sentence. BERT also uses a sentence-level embedding, known as , which can be used as a full-response representation rather than using distinct representations for each token. From here on, we will refer to this encoder as BERTfrozen. Additionally, can be fine-tuned by training BERT’s previously trained weights on new downstream classification tasks [49]. We will refer to this encoder as BERTfinetuned. While requiring more computational resources, this allows for the embedding to capture patterns in the text that are specific to the target prediction task.
@@
Shared classifiers across problems. Given a generated embedding for a student explanation, we want to classify the representation into its correct criteria vector . Thus, we have a multi-label task [50]. If we construct separate binary classification models for each criteria question, then we are required to train independent classifiers. Rather, we propose to work off of an informed assumption: shared domains will have similar word occurrence and grammatical structure across P problems. For example, the phrase “speeding up” should indicate evidence of sensemaking in the movement domain regardless of the problem it appears in , and words that should be classified as objects will share similar positions in the sentence across problems due to object-subject-verb rules. Thus, for problems, we can train 8 independent classifiers. We reduce from a scaling linear in to constant in . Note that we are proposing to train 8 classifiers to obtain annotations for student explanations across problems, where each problem has a different number of criteria questions per domain . Thus, to obtain an annotation for each criteria question within a domain, we add a criteria context embedding to each text embedding for a given student explanation. The updated embedding vector is denoted as . The index specifies the criteria question that we would like for the classifier to label (e.g. Does the student identify the object positioning before applied force (e.g. upwards)?). We note that there are multiple possible ways to encode this context. For example, rather than adding the vectors to obtain the updated embedding, one could concatenate the two vectors together. Additionally, one could embed the entire problem text alongside the specific criteria question. As a starting point, we’ve selected our approach of adding the embeddings together and only including the criteria question. We leave the exploration of alternative methods for future work.
@@
Figure 2 visually illustrates how the model makes a probabilistic prediction answering a specific yes/no criteria question given a student-generated text response. Given embedding , we use a domain-specific logistic regression (LR) classifier to produce the predicted probability . We set , where are the weight vector and bias scalar for domain , and denotes the logistic sigmoid function.
@@
The figure also shows how that prediction impacts the loss used to train the model. For training, we have a dataset student embeddings with human annotated labels (the criteria vectors) . Each represents a single binary label with respect to criteria question of domain for the student explanation . Each is assumed conditionally independent of other labels given and all model parameters.
We minimize the following binary cross entropy loss with an added regularization term over the weights averaged over all explanations, domains, and criteria:
@@
For the precise loss notation utilized in our methods including regularization terms, see Appendix C. The training procedure encourages all 8 classifiers to accurately predict their assigned pieces of the criteria vector for an embedded student explanation .
@@
For some encoder choices, we do not update any encoder parameters during training; only the classifier weights and biases are updated via gradient descent. When fine-tuning BERT, our gradient descent procedure updates encoder parameters alongside classifier weights. Both are explored in our deployability analysis in IV.1.
III.2 Theoretical Benefits and Limitations
Utility. The proposed model maps new student text explanations across a set of problems to sensemaking scores , freeing up time for an annotator. However, the current version of this model is limited to measuring sensemaking scores of student explanations on the same set of problems. We hypothesize that with more labeled training data across new problems, the model proposed here will have enough data to perform well on unseen problem instances in physics. However, the amount of data needed to cross such a threshold is unclear. Regardless, initial efforts to address this limitation should be directed towards collaborative annotation.
Reproducibility. The automation scheme ensures that the same student explanation receives the same annotation. The process is fully deterministic given fixed seeds to control for randomness. Randomness is only present in the initialization of the classifier weights and biases.
Sustainability. As ML research advances, new encoders may be able to improve the overall classification performance. However, many of these models require High Performance Computing (HPC) access and large amounts of energy resources for training [51]. As language modeling in ML improves, it is worth benchmarking new methods while also taking into account the amount of compute resources required to obtain only modestly better performance.
III.3 ML for Education Related Works
Machine learning tools are slowly integrating into STEM Ed. research methods, mostly with the introduction of text classification models [52, 43, 44, 45, 53]. One focus is on in-context learning, where student text responses are directly fed into a chatbot window of a large language model to be labeled [42]. This procedure can be conducted with a few examples (or zero) of student text, opening up an opportunity for non-ML practitioners to seamlessly use computation for data annotation. However, this procedure relies on the assumption that pre-trained language models are useful for specific tasks in education (e.g. classifying correctness, reasoning, sensemaking, etc.). Other works focus on training supervised ML models from scratch for their specific task [44, 45, 41]. For example, Watts et al. [41] uses the mechanistic reasoning approach from Russ et al. to analyze student explanations in organic chemistry with the aid of ML binary classifiers. This work determines convolutional neural networks to outperform other model architectures for this task, but does not provide comparison results to other models. In our work, we provide a model comparison with three different language encoders (BoW, BERTfrozen, and BERTfinetuned), where we also fine-tune a pre-trained large language model on data for our specific task. Furthermore, we use a multi-label scheme that allows us to take advantage of shared structure across problems by providing each classifier with additional context for the problem-specific criteria we care about. As such, our classifiers are able to train on and classify new data across different physics problems without introducing new classification heads.
In response to the increasing amount of supervised ML models in STEM Ed. research, Fussell et al. [54] effectively summarizes and provides a demonstration of the best practices widely used by ML practitioners for validating these ML models. We note that our work utilizes methods advised by this work, including both cross-validation and bootstrapping. Lastly, we note that while our work does not fall under the category of unsupervised learning, these models are a promising avenue of research in STEM Ed. and can be used to cluster student data, detect static latent variables in student text, and analyze the latent dynamics within student conversations [55].
IV Practical Implementation Results
IV.1 ML Model Deployability Analysis
@@
We seek to determine the deployability of the proposed classification models outlined in Sec. III.1. More specifically, our intention is to estimate how these models will perform given explanations from new students that it has never seen before, but on the same set of problems. Below, we outline our implementation and evaluation procedures prior to showcasing final results.
Dataset of student explanations. Our dataset consists of 385 total de-identified student explanations across four problems distributed to students in the first few lectures of the 2023 Tufts University Introduction to Physics course. The retrospective analysis of this deidentified human subjects data was approved by our institutional review board. All problems have the same structure as Problem 4 II.1 shown above and are presented to the students as pre-lecture digital assignments [38]. The same students are asked to solve each problem and explain how they obtained their solution. A few students chose not to respond to a particular problem, which brings our total number of explanations to . The problems capture different general core material for students to sensemake: (1) the descriptors of an object’s motion and dynamics; (2) the model of an object’s motion and dynamics under Earth’s gravitational influence; (3) the model of the relative motion of objects; (4) the model of an object’s motion dynamics under a net external influence. These problems vary in criteria vector size: (see Appendix A for exact problems and their details from [38]). The annotation procedure was carried out by a domain expert: the first author of this work who holds a doctorate in physics.
@@
Model and training specifics. In our analysis, we use the following models BoW, BERTfrozen, and BERTfinetuned. To evaluate performance, we randomly select 15 students who responded to all four questions to use as a held-out test set and use the rest of the data for training and validation (325 samples). To prevent data leakage, we stratify the data by students to ensure that responses from the same student do not end up in both training and test/validation sets.
@@
All models are trained with the Adam optimizer [56] for epochs. We perform a grid search over key model hyperparameters: learning rate and regularization strength ( for classifier weights and for BERT model weights), the full-details of which are available in Appendix D. For BoW and BERTfrozen, we use a 5-fold cross-validation approach for hyperparameter selection, where we select the hyperparameter combination with the best final average AUROC score across folds. AUROC is the computed area under a Receiver Operating Characteristic (ROC) curve defined by the true positive rate vs. the false positive rate of the predicted labels at different classification thresholds. There are two main reasons we believe AUROC to be the most useful metric for our context: (1) It does not require us to choose an arbitrary probability threshold to distinguish our positive predictions (label 1) from our negative predictions (label 0) for each classifier. Instead AUROC sums over all possible thresholds, providing a good summary that is still bounded, AUROC , with increasing values indicating better performance. (2) Due to imbalanced annotations (i.e. some classifiers have more of a positive or negative class rather than a 50-50 split) in our training dataset, AUROC is a safer metric to report than precision or accuracy, as the chance baseline is independent of imbalance. We include an annotation summary of the data in the Appendix B.

### v3

@@ III.2 Measurement as multi-label binary annotation
Step 1. Select problems from an introductory physics course on classical mechanics, where there exists a physical target phenomena to be understood.
@@
Step 2. For each problem and for each domain , determine the set of yes/no criteria questions that are pertinent to the specific problem. For example, in Problem 4 from Sec. II.1, there is only a single object, and thus with the criteria question Does the student identify the object?. However, there are two spatial organization criteria questions : Does the student identify the object spatial organization before the net applied force? Does the student identify the object spatial organization after the net applied force? We denote the total number of criteria questions for problem as . The full list of criteria for Problem 4 is displayed in the Appendix A.
@@
Step 3. Given a student’s text response to problem , the annotator will use the yes/no criteria questions to obtain a vector of binary digits representing their answers. The resulting criteria vector for a student response is denoted as . To further compute a sensemaking score one can sum the number of and divide by the length of the vector . This score is to be interpreted as the amount of evidence of mechanistic sensemaking. An annotator could decide to weight the domains (or specific criteria) differently. In this work, we keep the weights equal. To finish our example, the criteria vector for the student response in Fig. 1 is with the corresponding sensemaking score , a “yes” code for all domains but “property.” Considering the response itself, a high score seems appropriate: The student imagined a tangible scenario with multiple influences on the box.
III.3 Multi-label binary classification model
Here, we describe an ML model we constructed for our task, that can automatically map a student text response to a numerical sensemaking score, where a higher value indicates greater evidence of mechanistic sensemaking. While our approach relies on problem-specific criteria, some aspects of the model are shared across problems to gain statistical strength. The overall model is modular: components like the encoder can be varied to tradeoff between model capacity and resource consumption.
@@ III.3 Multi-label binary classification model
From text to embeddings. Given a student response , we can select an encoder to map all text in the response to a numerical representation . A simple encoder design is counting; we determine a vocabulary as the set of words across all student responses in the training dataset, with an additional vocabulary entry for unknown words, and record the number of times each one appears in the example. This numerical representation, or embedding , typically called Bag-of-Words, reduces the richness of the text explanation to the pattern of word occurrence [42, 43]. From here on, we will refer to this encoder as BoW.
@@
To capture more complex patterns (e.g relative co-occurrence between all words in the response, in-context word occurrence, and word positional associations with grammar rules), one can use a large language model (LLM) encoder [44]. For example, Google’s BERT (Bi-directional Encoder Representations from Transformers) model combines tokenization, self-attention, and a multi-layer perceptron neural network - all pre-trained on English Wikipedia and BookCorpus to generate token embeddings [1]. A token is a set of characters containing a unique ID in BERT’s vocabulary, such that an entire text can be converted to tokens with individual embeddings , which carry information about all other tokens in the passage. BERT also uses a sentence-level embedding, known as , which can be used as a full-response representation rather than using distinct representations for each token. From here on, we will refer to this encoder as BERTfrozen. As a final encoder option, the sentence embedding can be fine-tuned by training BERT’s previously trained weights on new downstream classification tasks [45]. We will refer to this encoder as BERTfinetuned. This allows for the embedding to capture patterns in the text that are specific to the target prediction task. It requires significantly more computational resources, however. Current LLMs, such as Google’s Gemini or OpenAI’s ChatGPT, require far more still. Part of our interest is to minimize the computational demands of the ML.
@@
Shared classifiers across problems. Given an embedding representation for a student response, we want to correctly assign it to a criteria vector . Thus, we have a multi-label task [46]. If we were to construct separate binary classification models for each criteria question, then we would be required to train independent classifiers. Rather, we propose to work off of an informed assumption: shared domains will have similar word occurrence and grammatical structure across problems. For example, the word “speeding” should indicate evidence of sensemaking in the movement activities domain regardless of the problem. Furthermore, words that should be classified as objects will share similar positions in the sentence across problems due to object-subject-verb rules. Thus, for problems, we propose to train 7 classifier heads, one for each domain. We reduce from a scaling linear in to constant in .
In order for this shared set of classifiers to effectively annotate multiple problems each with multiple yes/no criteria questions, we design each classifier to take an additional input beyond the embedding of student response text: a criteria context embedding that encodes the text of the yes/no question being asked (e.g. Does the student identify the object spatial organization before applied force (e.g. upwards)?). The index specifies the criteria question the classifier should label.
Each classifier must then take in student response embedding and criteria context embedding . There are multiple possible ways to combine the two inputs into a joint representation of response and context vectors that is the ultimate feature vector for the classifier. As a starting point, we have selected a simple approach of adding the embeddings together: . Possible alternatives would be to concatenate the two vectors together, or to produce one embedding from the combined text of question and response. We leave the exploration of alternative methods for future work.
@@
Fig. 2 visually illustrates how the model makes a probabilistic prediction for a specific yes/no criterion given a student’s response to a question. Given embedding , we use a domain-specific logistic regression (LR) classifier to produce the predicted probability . We set , where are the weight vector and bias scalar for domain , and denotes the logistic sigmoid function.
@@
The figure also shows how that prediction impacts the loss, which measures the difference between the model’s current predictions and human coding. For training, we have a dataset of student embeddings with human annotated labels (the criteria vectors) . Each represents a single binary label with respect to criteria question of domain for the student response . Each is assumed conditionally independent of other labels given and all model parameters. We minimize the following binary cross-entropy loss
@@
This loss is aggregated over all explanations , domains , and criteria . We also include an added regularization term that penalizes the sum-of-squares of the weight parameters. For the precise loss notation utilized in our methods including regularization terms, see Appendix C. The training procedure encourages all 7 classifiers to accurately predict their assigned pieces of the criteria vector for an embedded student response .
@@
For some encoder choices, we do not update any encoder parameters during training; only the classifier weights and biases are updated via gradient descent. When fine-tuning BERT, our gradient descent procedure updates encoder parameters alongside classifier weights. Both are explored in our model comparative analysis in Sec. IV.
Technical comparison with related work. Similar to Fussell et al. [8], we train a BagofWords (BoW) encoder and a BERTfinetuned encoder. We also observed that the BERTfinetuned encoder achieves better performance than the BoW encoder for our specific task. In general, our later results support their conclusion that fine-tuning adds value for a theory-driven classification in PER. They also compared two model architectures with multiple LLaMa encoders [2], which require significantly more computational resources than our BERT encoders. Similar to Watts et al. [28], we train multiple binary classifiers for task-specific evidence of mechanistic reasoning. They determined convolutional neural networks (CNNs) to outperform other model architectures such as naive Bayes, logistic regression, support vector machines, and transformers. Their results show high accuracy on the test data with slightly lower Cohen’s Kappa values to account for the data imbalance. We chose to continue with BERT encoders, expecting better inductive bias for our task than CNNs, especially when pretrained on large language corpora then fine-tuned on our limited data.
IV RQ2: How ML Design Choices Can Impact Performance
@@
We seek a performance comparison of the proposed classification models outlined in Sec. III.3 with different encoding and training strategies. More specifically, our intention is to estimate how these models will perform given explanations from new students that it has never seen before, but on the same set of problems. Below, we outline our implementation and evaluation procedures, then present and discuss results.
Dataset of student responses. Our dataset consists of 385 total de-identified student-responses across four problems distributed to 100 students in the first few lectures of the 2022 Tufts University Introduction to Physics course. The retrospective analysis of this deidentified human subjects data was approved by our institutional review board.
All problems have the same structure as Problem 4 in Sec. II.1 shown above and are presented to the students as pre-lecture digital assignments [23]. See Appendix A for exact problem text. The same students are asked to solve each problem and explain how they obtained their solution. A few students chose not to respond to a particular problem, which brings our total number of explanations to .
Annotating ground-truth labels. Complete annotation for all 4 problems was carried out by a domain expert: the first author of this work who holds a doctorate in physics. This included defining the yes/no criteria questions for each problem (provided in Appendix A) and assigning binary values for each student response. These problems vary in criteria vector size: . Some descriptive statistics summarizing annotations across problems are provided in Appendix B. We validated inter-rater reliability post-hoc, as described in Sec. V.3. In future work, we plan to incorporate training examples from multiple expert annotators.
@@
Experimental setup. To evaluate model performance on never-before-seen responses, we randomly select 15 students who responded to all four questions to use as a held-out test set. We use the rest of the data (325 samples) for training and validation. To prevent data leakage, we stratify the data by student to ensure that all responses from the same student across 4 problems end up either exclusively in the training set or exclusively in the test set.
Using this train/test split, we fit the models BoW, BERTfrozen, and BERTfinetuned described earlier. All models are trained with the Adam optimizer [47] for epochs. Convergence was verified by observing loss function behavior. We perform a grid search over key hyperparameters: learning rate and regularization strength ( for classifier weights and for BERT model weights); reproducible details are available in Appendix D. For BoW and BERTfrozen, we use 5-fold cross-validation, also stratified by student, to select hyperparameters. We pick the combination with the best average AUROC score across folds.
@@
The primary performance metric we report and evaluate is the AUROC, the Area Under the Receiver Operating Characteristic (ROC) curve. For a given classifier, this curve depicts the true positive rates (y-axis) and false positive rates (x-axis) achieved across different numerical thresholds for turning predicted probabilities into binary decisions. There are two main reasons we believe AUROC to be the most useful metric for our context: First, AUROC does not require us to choose an arbitrary probability threshold to distinguish label 1 predictions from label 0 predictions. Instead AUROC sums over all possible thresholds, providing an overall numerical summary where increasing values indicate better performance. AUROC is bounded in . Second, due to imbalanced annotations in our training dataset (i.e. some multi-label tasks have more of the positive or negative class rather than a 50-50 split), AUROC is an easier to interpret metric to report than precision or accuracy, as its chance baseline is independent of imbalance. For these reasons, we use AUROC as the primary performance metric throughout our analyses.
For the interested reader, we also include additional metrics like area under the precision-recall curve in the Appendix E and confusion matrices in Appendix F.